OpenStudy (anonymous):

The following data has been collected by a technician: Gas volume: V = 0.8 m3 Gas pressure: p = 160 kPa Compression index: n = 1.4 Pressure/volume relationship: pVn = C Work done = ∫_(V_1)^(V_2)▒〖p dV〗 You first need to work out C, the value of the pressure/volume relationship. Calculate the work done in compressing the nitrogen from 0.8 m3 to 0.6 m3 using: integral calculus Simpson’s rule

OpenStudy (anonymous):

this is simple ,just go for integration for given limits

OpenStudy (anonymous):

and for simpson's 3rd rule go for some spacing like from 0.8 to 0.6 in steps as from 0.8 to o.75 .then from 0.75 to .70 and so on then apply simpson"s formulae

OpenStudy (anonymous):

\int_{a}^{b} f(x) \, dx \approx \frac{b-a}{6}\left[f(a) + 4f\left(\frac{a+b}{2}\right)+f(b)\right].